# Quantum algorithms in distributed quantum computing

### v2

@@ Abstract
Distributed quantum computing (DQC), provides a way to scale quantum computers using multiple quantum processing units (QPU) which are connected through quantum communication links. In this paper, we have built a distributed quantum computing simulator and used the simulator to investigate quantum algorithms such as the quantum Fourier transform, quantum phase estimation, quantum amplitude estimation, and the generation of probability distribution in DQC. The simulator can be used to generate and execute distributed quantum circuits with ease, obtain and benchmark DQC parameters such as the fidelity of the algorithm and the number of entanglement generation steps, and use dynamic circuits in a distributed setting to improve results. We show the applicability of dynamic quantum circuits in DQC, where mid-circuit measurements, local operations, and classical communication are used in place of noisy inter-processor (nonlocal) quantum gates.
@@ 1 Introduction
To make quantum computers advantageous in industry, the scalability of the hardware needs to be significantly increased [Buhrman2003, Beals2013, Ferrari2021, Parekh2021]. Increasing the number of qubits increases the crosstalk, thereby increasing the gate and measurement error rates [PhysRevA.104.052404]. Distributed quantum computing (DQC) [Davis2023, Gyongyosi2021] offers an interesting approach. Instead of scaling up, we use several quantum processing units (QPUs) with fewer qubits connected through quantum communication links to increase computational capabilities [Lim2005, Niu2023]. The nodes of the DQC (Fig. 1) consist of the local qubits necessary for computation, and communication qubits [PhysRevLett.118.250502], which are used to entangle the nodes and apply quantum gates between them [Eisert2000, Monroe2009, Krutyanskiy2019].
@@
The communication qubits can be entangled using a probabilistic entanglement generation procedure that relies on successful detector click patterns [Nature2013, PhysRevLett.124.110501, PhysRevLett.120.200501]. The generation process is repeated until successful click patterns are obtained, indicating that the qubits have been entangled. Entanglement generation between remote nodes has been experimentally realized in superconducting qubits [PhysRevLett.120.200501], ions [PhysRevLett.124.110501], and NV centers [Nature2013] respectively. We will briefly study the success probability of this procedure.
Suppose the total efficiency (including qubit-node coupling) of the channel is $p$. For $k$ repetitions (trials), the probability of obtaining at least one entangled pair with 1 communication qubit/node is given by $P=(1-{(1-p)}^{k})$. To increase the probability of success, we can have multiple communication ($m$) qubits per node, where the entanglement generation procedure is run for each pair in parallel. The probability that we have at least one entangled pair is given by, $P_{s}=1-(1-P)^{m}$. The entangled pair is used to apply a remote quantum gate between local qubits in the different nodes utilising one ebit/CNOT gate by applying the quantum circuit shown in Fig. 2 [Eisert2000].
@@
Figure 2: The quantum circuit for the non-local teleportation-based CNOT gate between $q_{0}$ and $q_{3}$ consists of the following steps: a) The initial quantum state of $q_{1}$ and $q_{2}$ is prepared in the Bell state $\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle$), b) two local CNOT gates, followed by measurement on the qubits $q_{1}$ and $q_{2}$, classical communication and conditional single-qubit quantum gates
@@
The non-local CNOT gate has been experimentally realized in superconducting qubits [Chou2018] and ion qubits [Daiss2021]. A non-local quantum circuit could be obtained using multiple non-local quantum gates, where the gates are applied sequentially through entangling, measuring, and re-entangling the same qubit with only one extra qubit/node.
@@
In this paper, we describe a distributed quantum computing simulator (DQCS) that uses non-local quantum gates between the nodes, creates distributed quantum circuits, and evaluates their performance in the presence of noise [Eisert2000]. We use the DQCS to study distributed quantum algorithms [Zhou2023, Cirac1999, Yimsiriwattana2004], demonstrate the applicability of dynamic quantum circuits in DQC that reduce resource overhead, and achieve high fidelity. We illustrate this with an example of distributed quantum phase estimation (DQPE). We study the probability distribution loading in DQC, scaling with the number of nodes. We use the DQCS to run the distributed quantum amplitude estimation (DQAE) algorithm on multiple nodes. The DQCS contains modules that provide users various ways to study, run, and evaluate the performance of quantum algorithms in DQC [Cirac1999]. The DQCS could be employed for simulating and optimizing both the quantum network layer and the quantum circuit layer for distributed quantum algorithms.
Several distributed quantum computing [Ferrari2021, Rhea2021, Hanner2021, Cuomo2020] and quantum internet simulators [Dahlberg_2019, Dahlberg_2022, QUISP] that have been developed that consider distributed quantum gates, compilers, and classical networks respectively. The goal of the DQCS is not only to consider the realistic setting for DQC, including node efficiency, gate imperfections, and dynamic quantum circuits to improve fidelity, which are crucial for experiments, but also to investigate and benchmark the performance of quantum algorithms in DQC.
@@ 2 Noise model for non-local quantum gates
The noise model for non-local quantum gates can be studied using the depolarization channel $E$ that acts on the quantum state of $n$ qubits $\rho$, $\rho\rightarrow E(\rho)$ as,
@@
where $\epsilon_{d}$ is considered as the depolarization single-qubit gate error rate (respectively) and $Tr_{i}$ is the partial trace of the $i^{th}$ qubit. Let us consider a two-qubit quantum gate $U_{i,j}$ (between the qubits $i$ and $j$). The two-qubit quantum gate error channel could be obtained for $n=2$ that applies the transformation $U_{ij}\rho U_{ij}^{\dagger}\rightarrow\rho^{\prime\prime}$.
@@ 3 Distributed quantum computing simulator
  • Distributed circuits: The distributed circuits module enables users to effortlessly create distributed quantum circuits using a node dictionary and a quantum circuit. It automatically handles communication registers, so users don’t need to explicitly specify them.
@@
  • Noise model: We can experiment with the distributed quantum circuit in the presence of noise, the library offers an optional noise model module. This component incorporates realistic noise modeling, including single-qubit and two-qubit errors. The noise parameters can be changed corresponding to different experimental systems and the fidelity of the quantum algorithm can be obtained. Note that in place of the entanglement generation, the Bell pair is created using a CNOT gate in the simulator. Hence, there is a small difference in the noise model where the depolarization of the Bell pair is approximated with the gate error $\epsilon_{g}/2$.
@@
    The circuit gives the distributed quantum circuit the parameter $gate\textunderscore app=1$ upon the successful creation of the distributed quantum circuit.
@@
    Instantiate the noise model with your distributed quantum circuit, specifying parameters such as the number of shots, repetitions, and error probabilities (single, CNOT quantum gates, measurement, reset errors). We show an example of including a noisy $CNOT$ gate in the simulation.
@@ 4 Distributed quantum algorithms
Quantum algorithms are utilized to achieve speedup for various applications including machine learning, optimization [Farhi2014], credit risk analysis [Egger2021], and option pricing using Monte Carlo integration [Nik2020, Carrera2021, Herbert2022]. We consider the simulation of two distributed quantum algorithms, DQPE and DQAE, in DQC which are the underlying quantum algorithms for more advanced quantum algorithms [Egger2021, Nik2020, Carrera2021, Herbert2022].
@@ 4.1 Dynamic quantum circuits
Dynamic quantum circuits involve the use of mid-circuit measurements, qubit reset, reuse, and single-qubit conditional quantum gates [PhysRevLett.127.100501]. Here, we explore the use of dynamic quantum circuits instead of distributed quantum gates in the DQC [dynamic]. One of the examples of dynamic quantum circuits is when we have a quantum Fourier transform followed by measurement [dynamic].
@@
The corresponding dynamic quantum circuit is shown in Fig. 4.
@@
We can now illustrate the application of distributed dynamic QFT to the quantum phase estimation.
@@ 4.2 Distributed dynamic quantum phase estimation
The problem of quantum phase estimation is to determine the eigenvalue of the unitary gate that acts as $U|\psi\rangle=e^{2i\pi y}|\psi\rangle$, where $|\psi\rangle$ is the eigenstate of $U$, and $y$ is a fraction corresponding to the eigenvalue. The quantum circuit used to implement QPE is shown in Fig. 5.
@@
We can derive the fraction $y$ from the outcome of the measurement. For instance, if we obtain the measurement outcome $y_{0}y_{1}y_{2}$, where each $y_{i}\in{0,1}$, then the fraction $y$ is calculated as $(y_{0}y_{1}y_{2})_{10}/2^{3}$. To illustrate, let’s examine the implementation of Quantum Phase Estimation (QPE) for the unitary gate $U=R_{X}(\pi)\otimes R_{Y}(\pi)$ acting on two qubits. One of the eigenvalues for this operator is $-1$ ($y=1/2$), corresponding to a measurement output of $100$. We prepare the eigenstate $|\psi\rangle$, create the distributed QPE (DQPE) for the following qubit distributions: 1: [0, 1, 2], 2: [3, 4], and apply the dynamic QFT to the distributed quantum circuit. The probability of the outcome is shown in Fig. 6, where we consider the locally controlled unitary gates and quantum Fourier transform for the QPE (dark blue) (Fig. 3). For the DQPE (blue), we consider distributed two-qubit quantum gates and dynamic QFT without two-qubit quantum gates for the quantum Fourier transform (Fig. 4). From the noise model, the dynamic quantum Fourier transform doesn’t suffer from two-qubit quantum gate errors, only measurement and depolarization errors, giving better results in Fig. 6. For quantum phase estimation using $n$ control qubits, we need $N$ two-qubit quantum gates to realize the controlled-unitary operations and the quantum Fourier transform requires $n(n-1)/2$ two-qubit quantum gates. Assuming the same error for the two-qubit (controlled-unitary and controlled phase) gates, the fidelity of the algorithm is approximated $F\approx(1-\epsilon_{g})^{N+\frac{n}{2}(n-1)}$. For the DQPE, where we have $N$ distributed quantum gates (each applied non-locally using 2 CNOT quantum gates), the fidelity is given by $F\approx(1-\epsilon_{g})^{2N}$.
@@
Figure 6: The probability of the outcome for the distributed QPE (DQPE) (blue) vs. QPE (red) in the presence of noise $\epsilon_{d}=\epsilon_{g}=\epsilon$, $\epsilon_{m}=0$

### v3

# The simulation of distributed quantum algorithms

@@ Abstract
Distributed quantum computing (DQC) provides a way to scale quantum computers using multiple quantum processing units (QPU) connected through quantum communication links. In this paper, we have built a distributed quantum computing simulator and used the simulator to investigate quantum algorithms such as the quantum Fourier transform, quantum phase estimation, quantum amplitude estimation, and generation of probability distribution in DQC. The simulator can be used to easily generate and execute distributed quantum circuits, obtain and benchmark DQC parameters such as the fidelity of the algorithm and the number of entanglement generation steps, and use dynamic circuits in a distributed setting to improve results. We show the applicability of dynamic quantum circuits in DQC, where mid-circuit measurements, local operations, and classical communication are used in place of noisy inter-processor (non-local) quantum gates.
@@ 1 Introduction
To make quantum computers advantageous in industry, the scalability of the hardware needs to be significantly increased [Buhrman2003, Beals2013, Ferrari2021, Parekh2021]. Due to the hardware limitations, increasing the qubits in the same device, increases the cross talk, decoherence that causes the effective gate and the measurement errors to increase [Nicholas2025]. Distributed quantum computing (DQC) [Davis2023, Gyongyosi2021] offers an interesting approach, where instead of scaling up, we use multiple quantum processing units (QPUs) with fewer qubits connected through quantum communication links to increase computational capabilities [Lim2005, Niu2023]. The QPUs can be separated by short distances, i.e., chiplets connected through microwave links [Nicholas2025] or long distances connected through optical fibers/photonic links [Main2025]. The nodes of the DQC (Fig. 1) consist of the local qubits necessary for computation and communication qubits [PhysRevLett.118.250502], which are used to entangle the nodes and apply quantum gates between them [Eisert2000, Monroe2009, Krutyanskiy2019, Main2025].
@@
The communication qubits can be entangled using a probabilistic entanglement generation procedure that relies on successful detector click patterns [Nature2013, PhysRevLett.124.110501, PhysRevLett.120.200501]. The generation process is repeated until successful click patterns are obtained, indicating that the qubits are entangled. Entanglement generation between remote nodes has been experimentally realized in superconducting qubits [PhysRevLett.120.200501], ions [PhysRevLett.124.110501, Main2025], and NV centers [Nature2013] respectively. We will briefly study the probability of success of this procedure.
Suppose the channel’s total efficiency (including qubit-node coupling) is $p$. For $k$ repetitions (trials), the probability of obtaining at least one entangled pair with one communication qubit/node is given by $P=(1-{(1-p)}^{k})$. To increase the probability of success, we can have multiple communication ($m$) qubits per node, where the entanglement generation procedure is run for each pair in parallel. The probability that we have at least one entangled pair is given by $P_{s}=1-(1-P)^{m}$. The entangled pair is used to apply a remote quantum gate between local qubits in the different nodes utilizing one ebit/CNOT gate by applying the quantum circuit shown in Fig. 2 [Eisert2000].
@@
Figure 2: The quantum circuit for the non-local teleportation-based CNOT gate between $q_{0}$ and $q_{3}$ using the Bell pair $q_{1}$ and $q_{2}$ prepared in $\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle$), followed by two local CNOT gates, measurement on the qubits $q_{1}$ and $q_{2}$, classical communication and conditional single-qubit quantum gates
@@
The non-local CNOT gate has been experimentally realized in superconducting qubits [Chou2018] and ion qubits [Daiss2021]. A non-local quantum circuit could be obtained using multiple non-local quantum gates, where the gates are applied sequentially through entangling, measuring, and re-entangling the same qubit with only one extra qubit/node. Recently, distributed Grover’s algorithm has been experimentally realized using dual species trapped ions connected through photonic link, where the nodes were separated by a distance of $2$ meters [Main2025].
@@
In this paper, we describe a distributed quantum computing simulator (DQCS) that uses non-local quantum gates between the nodes, creates distributed quantum circuits, and evaluates their performance in the presence of noise [Eisert2000]. We use DQCS to investigate and benchmark the performance of several distributed quantum algorithms [Zhou2023, Cirac1999, Yimsiriwattana2004], demonstrating the applicability of dynamic quantum circuits in DQC that reduce resource overhead and achieve high fidelity. We illustrate this with a distributed quantum phase estimation (DQPE) example. We study the loading of the probability distribution in DQC, scaling with the number of nodes. We use DQCS to run the distributed quantum amplitude estimation (DQAE) algorithm on multiple nodes. The DQCS contains modules that provide users various ways to study, run, and evaluate the performance of quantum algorithms in DQC [Cirac1999]. The DQCS could simulate and optimize the quantum network and circuit layers for distributed quantum algorithms.
Several distributed quantum computing [Ferrari2021, Rhea2021, Hanner2021, Cuomo2020] and quantum internet simulators [Dahlberg_2019, Dahlberg_2022, QUISP] that have been developed that consider distributed quantum gates, compilers, and classical networks respectively. They consider the control platform for the hardware (Interlin-q) [Rhea2021], job schedulers [Ferrari2021], optimization using the quantum message parsing interface (QMPI) [Hanner2021] for the simulation of DQC, analogous to the high performance computing. They consider the noiseless setting that does not include the gate, measurement imperfections, coupling efficiency, etc. that is crucial to realize DQC in the Noisy intermediate-scale quantum era (NISQ) quantum devices. In addition, they don’t consider other techniques, such as dynamic quantum circuits in the distributed setting, that reduce the number of non-local quantum gates. The key challenge of the DQC is that each non-local two-qubit quantum gate requires two local two-qubit gates, two measurements, and efficient coupling between the nodes. Thus, for a given set of hardware parameters (such as gate fidelity, measurement error rate, and coupling efficiency), it’s crucial to study the performance of a quantum algorithm [PhysRevA.104.052404] in the DQC in the presence of various imperfections. The goal of the DQCS is to consider the implementation of DQC in the NISQ devices [Main2025], consider the hardware parameters including node efficiency (for entanglement generation), gate imperfections (for non-local quantum gates), and dynamic quantum circuits to improve fidelity, which are crucial for experiments. Further, since the DQCS is designed for the gate-model quantum computation, established techniques, such as circuit depth reduction [Gyongyosi2020], are crucial and can be applied to improve the time complexity of the distributed quantum algorithm. Additionally, the latency introduced by classical communication in the DQC can be minimized by optimizing qubit distribution [Martinez2023], utilizing routing protocols [Hahn2019, Kristjánsson2024, Gyongyosi2022], different protocols for entanglement generation [Gyongyosi2022, Muralidharan2016] and entanglement swapping to maximize the rate, enhance the fidelity of the quantum algorithm [Gyongyosi2018].
The paper is organized as follows: Section 2 introduces the noise model for non-local quantum gates. Section 3 introduces the DQCS using code snippets that enable us to obtain the distributed quantum circuit using the DQCS package. In sections 4.1, 4.2, we introduce dynamic quantum circuits for distributed quantum phase estimation and show that the advantage improves as we increase the number of qubits. In section 4.3, we estimate the resources for the distributed quantum amplitude estimation algorithm for uniform distribution and estimate the fidelity for the probability distribution loading of normal distribution for different numbers of nodes in the DQC.
@@ 2 Noise model for non-local quantum gates
The noise model for non-local quantum gates can be studied using the depolarization channel $E$ that acts on the quantum state of $n$ qubits $\rho$, $\rho\rightarrow E(\rho)$ as
@@
where $\epsilon_{d}$ is considered as the depolarization single-qubit gate error rate (respectively) and $Tr_{i}$ is the partial trace of the $i^{th}$ qubit. Consider a two-qubit quantum gate $U_{i,j}$ (between the qubits $i$ and $j$). The two-qubit quantum gate error channel could be obtained for $n=2$ that applies the transformation $U_{ij}\rho U_{ij}^{\dagger}\rightarrow\rho^{\prime\prime}$.
@@ 3 Distributed quantum computing simulator
  • Distributed circuits: The distributed circuits module enables users to create distributed quantum circuits using a node dictionary and a quantum circuit. It automatically handles communication registers, so users don’t need to specify them explicitly.
@@
  • Noise model: We can experiment with the distributed quantum circuit in the presence of noise; the library offers an optional noise model module. This component incorporates realistic noise modeling, including single-qubit and two-qubit errors. The noise parameters can be changed to correspond to different experimental systems, and the fidelity of the quantum algorithm can be obtained. Note that the Bell pair is created using a CNOT gate in the simulator instead of the entanglement generation. Hence, there is a small difference in the noise model where the depolarization of the Bell pair is approximated with the gate error $\epsilon_{g}/2$.
@@
    The circuit gives the distributed quantum circuit the parameter $gate\textunderscore app=1$ upon successfully creating the distributed quantum circuit.
@@
    Instantiate the noise model with your distributed quantum circuit, specifying the number of shots, repetitions, and error probabilities (single, CNOT quantum gates, measurement, reset errors). We show an example of including a noisy $CNOT$ gate in the simulation.
@@ 4 Distributed quantum algorithms
Quantum algorithms are utilized to achieve speedup for various applications, including machine learning, optimization [Farhi2014], credit risk analysis [Egger2021], and option pricing using Monte Carlo integration [Nik2020, Carrera2021, Herbert2022]. We consider the simulation of two distributed quantum algorithms, DQPE and DQAE, in DQC, which are the underlying quantum algorithms for more advanced quantum algorithms [Egger2021, Nik2020, Carrera2021, Herbert2022].
@@ 4.1 Dynamic quantum circuits
Dynamic quantum circuits involve using mid-circuit measurements, qubit reset, reuse, and single-qubit conditional quantum gates [PhysRevLett.127.100501]. Here, we explore using dynamic quantum circuits instead of distributed quantum gates in the DQC [dynamic]. One of the examples of dynamic quantum circuits is when we have a quantum Fourier transform followed by measurement [dynamic].
@@
Fig. 4 shows the corresponding dynamic quantum circuit.
@@
Note that the dynamic QFT does not suffer from two-qubit gate errors; it suffers from depolarization errors $\epsilon_{d}=1-e^{-t/T_{2}}$ [Piparo2024], where t is the classical communication time, and $T_{2}$ is the coherence time of the qubit. For $T_{2}=50\mu s$, QPUs separated by $10m$, speed of light in fiber $c=2\times 10^{8}m/s$, we get $\epsilon_{d}\approx 10^{-3}$. We can now illustrate the application of distributed dynamic QFT to the quantum phase estimation.
@@ 4.2 Distributed dynamic quantum phase estimation
The problem of quantum phase estimation (QPE) is to determine the eigenvalue of the unitary gate that acts as $U|\psi\rangle=e^{2i\pi y}|\psi\rangle$, where $|\psi\rangle$ is the eigenstate of $U$, and $y$ is a fraction corresponding to the eigenvalue. The quantum circuit used to implement QPE is shown in Fig. 5.
@@
We can derive the fraction $y$ from the outcome of the measurement. For instance, if we obtain the measurement outcome $y_{0}y_{1}y_{2}$, where each $y_{i}\in{0,1}$, then the fraction $y$ is calculated as $(y_{0}y_{1}y_{2})_{10}/2^{3}$. To illustrate, let’s examine the implementation of QPE for the unitary gate $U=R_{X}(\pi)\otimes R_{Y}(\pi)$ acting on two qubits. One of the eigenvalues for this operator is $-1$ ($y=1/2$), corresponding to a measurement output of "$100"$. We prepare the eigenstate $|\psi\rangle$, create the distributed QPE (DQPE) for the following qubit distributions: 1: [0, 1, 2], 2: [3, 4], and apply the dynamic QFT to the distributed quantum circuit. The probability of the measurement outcome "$100"$ is shown in Fig. 6, where we consider the locally controlled unitary gates and quantum Fourier transform for the QPE (red) (Fig. 3). For the DQPE (blue), we consider distributed two-qubit quantum gates and dynamic QFT without two-qubit quantum gates for the quantum Fourier transform (Fig. 4). The probability of measuring "$100"$ is shown in Fig. 6. For the DQPE (blue), we consider distributed two-qubit quantum gates and dynamic QFT without two-qubit quantum gates for the quantum Fourier transform (Fig. 4). We consider the gate error $\epsilon_{g}$, depolarization error $\epsilon_{d}$, and the measurement error $\epsilon_{m}$. In Fig. 6, we consider the following scenarios
  1. 1.
    Probability vs $\epsilon_{g}$, where $\epsilon_{g}=10\epsilon_{d}$, $\epsilon_{m}=0$ (Fig. 6(a)).
  2. 2.
    Probability vs $\epsilon_{g}$, where $\epsilon_{g}=10\epsilon_{d}$, $\epsilon_{m}=\epsilon_{g}/4$ (Fig. 6(b)).
While each distributed quantum gate requires two local CNOT gates, the dynamic quantum Fourier transform doesn’t suffer from two-qubit quantum gate errors, only measurement and depolarization errors, giving better results in Fig. 6. Since each non-local quantum gate requires measurement that is not present in the local implementation of a quantum circuit, we consider the measurement error in our analysis. Note that there are efficient methods to reduce the measurement error to $\epsilon_{g}/4$ [Knill2005, Muralidharan2016].
(a) Probability vs $\epsilon_{g}$, $\epsilon_{d}=\epsilon_{g}/10$, $\epsilon_{m}=0$
(b) Probability vs $\epsilon_{g}$, $\epsilon_{d}=\epsilon_{g}/10$, $\epsilon_{m}=\epsilon_{g}/4$
@@
Figure 6: The probability measuring the required bit for the distributed QPE (DQPE) (blue) vs. QPE (red) in the presence of depolarization, gate, measurement errors $\epsilon_{d}$, $\epsilon_{g}$, $\epsilon_{m}$ respectively.
The improvement obtained with DQPE increases when considering a larger number of qubits. Let us consider the inverse QFT on five qubits (Fig. 7) that uses ten two-qubit and five Hadamard gates.
[…]
Figure 7: Quantum Fourier transform (inverse) on five qubits, where $P(\theta)$ refers to the phase gate
The dynamic quantum circuit is given in Fig. 8,
[…]
Figure 8: Dynamic quantum Fourier transform using five qubits
The quantum circuit used to implement the DQPE is shown in Fig. 9.
[…]
Figure 9: Distributed dynamic quantum phase estimation for five qubits where the qubits are in different nodes. The controlled unitary gates are applied using DQCS, followed by dynamic QFT
We estimate the fraction $y$ that is calculated as $(y_{0}y_{1}y_{2}y_{3}y_{4})_{10}/2^{5}$. We prepare the eigenstate $|\psi\rangle$, create the distributed QPE (DQPE) for the following qubit distributions: 1: [0, 1, 2, 3, 4], 2: [5, 6], and apply the dynamic QFT to the distributed quantum circuit. We consider the following scenarios:
  1. 1.
    Probability of obtaining the correct outcome vs error rate $\epsilon$, where $\epsilon_{g}=\epsilon_{d}=\epsilon$, $\epsilon_{m}=0$ (Fig. 10(a)).
  2. 2.
    Probability vs the two-qubit gate error $\epsilon_{g}$, where $\epsilon_{g}=10\epsilon_{d}$, $\epsilon_{m}=0$ (Fig. 10(b)).
  3. 3.
    Probability vs $\epsilon$, where $\epsilon_{g}=\epsilon_{d}=\epsilon$, $\epsilon_{m}=\epsilon_{g}/4$ (Fig. 10(c)).
  4. 4.
    Probability vs $\epsilon_{g}$, where $\epsilon_{g}=10\epsilon_{d}$, $\epsilon_{m}=\epsilon_{g}/4$ (Fig. 10(d)).
In Fig. 10, we note that there is better improvement obtained using DQPE (using five qubits) than three qubits (Fig. 6). This could be understood theoretically as follows: For quantum phase estimation using $n$ control qubits, suppose that we need $N$ two-qubit quantum gates to realize the controlled-unitary operations, and the quantum Fourier transform requires $n(n-1)/2$ two-qubit quantum gates. Theoretically, assuming the same error for the two-qubit (controlled-unitary and controlled phase) gates, the fidelity of the algorithm is approximated $F\approx(1-\epsilon_{g})^{N+\frac{n}{2}(n-1)}$. For DQPE, where we have $N$ distributed quantum gates (each applied using two CNOT quantum gates), the fidelity is given by $F\approx(1-\epsilon_{g})^{2N}$.
(a) $\epsilon_{g}=\epsilon_{d}=\epsilon$, $\epsilon_{m}=0$
(b) $\epsilon_{d}=\epsilon_{g}/10$, $\epsilon_{m}=0$
(c) $\epsilon_{g}=\epsilon_{d}=\epsilon$, $\epsilon_{m}=\epsilon_{g}/4$
(d) $\epsilon_{d}=\epsilon_{g}/10$, $\epsilon_{m}=\epsilon_{g}/4$
Figure 10: The probability of the correct outcome for the distributed QPE (DQPE) (blue) vs. QPE (red) in the presence of gate, depolarization, and measurement errors